# A Survey on Large Language Model based Human-Agent Systems

### v1

@@ A Survey on Large Language Model based Human-Agent Systems
Henry Peng Zou1, *, Wei-Chieh Huang1, *, Yaozu Wu2, *, Yankai Chen1, †, Chunyu Miao1, Hoang Nguyen1, Yue Zhou1, Weizhi Zhang1, Liancheng Fang1, Langzhou He1, Yangning Li3, Yuwei Cao4, Dongyuan Li2, Renhe Jiang2, Philip S. Yu1, † 1University of Illinois Chicago, 2The University of Tokyo, 3Tsinghua University, 4Google DeepMind {pzou3, whuang80, psyu}@uic.edu, yaozuwu279@gmail.com, yankaichen@acm.org
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have led to growing enthusiasm for building fully autonomous agents — systems that use LLMs as a central engine to perceive environments, make decisions, and execute actions to achieve goals Wang et al. (2024a); Li et al. (2024a). These agents are often equipped with modules for memory, planning, and tool use, aiming to automate complex workflows with minimal human involvement Xi et al. (2025). However, the pursuit of full autonomy faces critical hurdles. (1) Reliability remains a major concern due to LLMs’ propensity for hallucination – generating plausible but factually incorrect or nonsensical outputs – which erodes trust and can lead to significant errors, especially when actions are chained Gosmar and Dahl (2025); Xu et al. (2024). (2) Complexity often stalls autonomous agents; they struggle with very complicated tasks requiring deep domain expertise, long multi-step execution, nuanced reasoning, dynamic adaptation, or strict long-context consistency dependencies, as seen in scientific research Feng et al. (2024); Yehudai et al. (2025). (3) Safety and Ethical Risks escalate with autonomy; agents can take unintended harmful actions, amplify societal biases present in training data, or create accountability gaps, particularly in critical decision-making scenarios involving finance, healthcare, or security Mitchell et al. (2025); Deng et al. (2024); Shen et al. (2024).
@@
Figure 1: Overview of LLM-based human-agent systems. The system is composed of five core components: Environment & Profiling (including environment settings, and role definitions, goals, and agent capabilities such as planning and memory), Human Feedback (with varying types, timing, and granularity), Interaction Types (collaborative, competitive, cooperative, or mixed), Orchestration (task strategy and temporal synchronization), and Communication (information flow structure and mode). Together, these five components define the structure and functionality of LLM-based human–agent systems.
@@
The persistence of these challenges suggests that full autonomy may be unsuitable for many real-world applications Mitchell et al. (2025); Natarajan et al. (2025) and underscores a crucial insight often overlooked in the drive for pure automation: the indispensable role of human involvement. Humans are frequently needed to provide essential clarification, context, or domain knowledge, offer vital feedback and corrections, and exercise necessary oversight and control. These motivate a paradigm shift towards systems explicitly designed for human-agent collaboration: LLM-based Human-Agent Systems (LLM-HAS).
@@
While surveys on LLM-based autonomous agents Wang et al. (2024a); Li et al. (2024a), multi-agent systems Tran et al. (2025); Wu et al. (2025), and specific applications exist Wang et al. (2025b); Peng et al. (2025), a dedicated synthesis focusing specifically on LLM-based human-agent systems is lacking. This survey fills that gap by providing a comprehensive and structured overview of the LLM-HAS. It clarifies the fundamental concepts and systematically presents its core components, emerging applications, and unique challenges and opportunities within this specific niche. To the best of our knowledge, this is still the first survey on LLM-based human-agent systems. We aim to consolidate current knowledge and inspire further research and innovation in this rapidly evolving interdisciplinary field.
To provide a sustainable resource complementing our survey paper, we maintain an open-source GitHub repository. We hope that our survey will inspire further exploration and innovation in this field, as well as applications across a wide array of research disciplines.
@@
This survey is organized as follows: Section 2 defines and formulates LLM-HAS. Section 3 details the core components shaping the human-agent systems (e.g., human feedback, interaction type, orchestration and communication protocols). Section 4 explores diverse application domains. Section 5 presents open-source implementation frameworks as well as datasets and benchmarks. Finally, Section 6 discusses key challenges and future opportunities in the LLM-based human-agent systems.
@@
Type, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Evaluative, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ CoELA Zhang et al. (2024a), DigiRL Bai et al. (2024),
Auto Agent Pan et al. (2024b), WebCanvas Pan et al. (2024d), MineWorld Guo et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Corrective, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Collaborative Gym Shao et al. (2024), MTOM Zhang et al. (2024b),
AI Chains Wu et al. (2022), HRC DMP Liu et al. (2024), MindAgent Gong et al. (2023), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Guidance, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ FineArena Xu et al. (2025), InteractGen Sun et al. (2024b), HRC Assembly Gkournelos et al. (2024),
ConvCodeWorld Han et al. (2025), DPT Agent Zhang et al. (2025), Organized Teams Guo et al. (2024b), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Implicit, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), REVECA Seo et al. (2025), AssistantX Sun et al. (2024a),
@@
ReHAC Feng et al. (2024), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Feedback
Granularity, fill= HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Holistic, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Help Feedback Mehta et al. (2024), PARTNR Chang et al. (2024), Auto Agent Pan et al. (2024b),
MINT Wang et al. (2024c), HRC DMP Liu et al. (2024), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font =] ] [ Segment, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ CowPilot Huq et al. (2025), SWEET-RL Zhou et al. (2025), PDFChatAnnotator Tang et al. (2024), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Feedback
Phase, fill= HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Initial Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ AgentCoord Pan et al. (2024a), AssistantX Sun et al. (2024a), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ During Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ ConvCodeWorld Han et al. (2025), HRC Manipulation Liu et al. (2023a), A2C Tariq et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Post-Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ CoELA Zhang et al. (2024a), Auto Agent Pan et al. (2024b),
@@
WebCanvas Pan et al. (2024d), MineWorld Guo et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] ] [ Interaction
@@
(Sec 3.3), fill=interactClr, text width= 7em, anchor=center, align=center, font=[ Collaboration, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Supervision, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Collaborative Gym Shao et al. (2024), CowPilot Huq et al. (2025), Hierarchical Agent Liu et al. (2023b),
ConvCodeWorld Han et al. (2025), HRC Manipulation Liu et al. (2023a), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Delegation, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Organized Teams Guo et al. (2024b), GDfC Wang et al. (2025c), HRC Manufa Lim et al. (2024),
PaLM-E Driess et al. (2023), DigiRL Bai et al. (2024), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Coordination, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ MTOM Zhang et al. (2024b), Ask Before Plan Zhang et al. (2024c), MINT Wang et al. (2024c),
AssistantX Sun et al. (2024a), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Cooperation, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ CoELA Zhang et al. (2024a), InteractGen Sun et al. (2024b), PARTNR Chang et al. (2024),
Agency Task Sharma et al. (2024), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] ] [ Competition, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ SOTOPIA Zhou et al. (2024), leaf, fill=interactLight, text width=51.6em, anchor=center, align=center, text centered, font = ] ] [ Coopetition, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), SOTOPIA Zhou et al. (2024), leaf, fill=interactLight, text width=51.6em, anchor=center, align=center, text centered, font = ] ] ] [ Orchestration
@@
Help Feedback Mehta et al. (2024), Ask Before Plan Zhang et al. (2024c), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] [ Simultaneous, fill=orchestLight, text width= 6.5em, anchor=center, align=center, font=[ MTOM Zhang et al. (2024b), Hierarchical Agent Liu et al. (2023b), DPT Agent Zhang et al. (2025),
@@
HA Comm Bansal et al. (2024), HRC Manufa Lim et al. (2024), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] [ Asynchronous, fill=orchestLight, text width= 6.5em, anchor=center, align=center, font=[ Auto Agent Pan et al. (2024b), DigiRL Bai et al. (2024), TaPA Wu et al. (2023), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] ] ] [ Communication
(Sec 3.5), fill=communClr, text width= 7em, anchor=center, align=center, font=[ Structure, fill=communLight, text width= 5em, anchor=center, align=center, font=[ Centralized, fill=communLight, text width= 6em, anchor=center, align=center, font=[ HRC Assembly Gkournelos et al. (2024), Drive As You Speack Cui et al. (2024),
MindAgent Gong et al. (2023), Ask Before Plan Zhang et al. (2024c), SOTOPIA Zhou et al. (2024), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Decentralized, fill=communLight, text width= 6em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), InteractGen Sun et al. (2024b), SWEET-RL Zhou et al. (2025),
MINT Wang et al. (2024c), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Hierarchical, fill=communLight, text width= 6em, anchor=center, align=center, font=[ FineArena Xu et al. (2025), AI Chains Wu et al. (2022),
EasyLAN Pan et al. (2024c), Hierarchical Agent Liu et al. (2023b), REVECA Seo et al. (2025), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Mode, fill=communLight, text width= 5em, anchor=center, align=center, font=[ Conversation, fill=communLight, text width= 6em, anchor=center, align=center, font=[ DigiRL Bai et al. (2024), WebLINX Lù et al. (2024), Organized Teams Guo et al. (2024b),
ConvCodeWorld Han et al. (2025), Help Feedback Mehta et al. (2024), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Observation, fill=communLight, text width= 6em, anchor=center, align=center, font=[ REVECA Seo et al. (2025), A ttentive Supp Tanneberg et al. (2024a), EasyLAN Pan et al. (2024c),
DPT Agent Zhang et al. (2025), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Message
Pool, fill=communLight, text width= 6em, anchor=center, align=center, font=[ InteractGen Sun et al. (2024b), AssistantX Sun et al. (2024a), MetaGPT Hong et al. (2023), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] ] ]
Figure 2: Taxonomy of LLM-based Human-Agent Systems.

### v2

@@ A Survey on Large Language Model based Human-Agent Systems
Henry Peng Zou1, *, Wei-Chieh Huang1, *, Yaozu Wu2, *, Yankai Chen1, †, Chunyu Miao1, Hoang Nguyen1, Yue Zhou1, Weizhi Zhang1, Liancheng Fang1, Langzhou He1, Yangning Li3, Dongyuan Li2, Renhe Jiang2, Xue Liu4,5, Philip S. Yu1, † 1University of Illinois Chicago, 2University of Tokyo, 3Tsinghua University, 4MBZUAI, 5McGill University {pzou3, whuang80, psyu}@uic.edu, yaozuwu279@gmail.com, yankaichen@acm.org
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have led to growing enthusiasm for building fully autonomous agent systems that use LLMs as a central engine to perceive environments, make decisions, and execute actions to achieve goals Wang et al. (2024a); Li et al. (2024a). These agents are often equipped with modules for memory, planning, and tool use, aiming to automate complex workflows with minimal human involvement Xie et al. (2024); Xi et al. (2025). However, the pursuit of full autonomy faces critical hurdles. (1) Reliability remains a major concern due to LLMs’ propensity for hallucination, generating plausible but factually incorrect or nonsensical outputs, which undermines trust and can lead to significant errors, especially when actions are chained Gosmar and Dahl (2025); Xu et al. (2024); Glickman and Sharot (2025). (2) Complexity often stalls autonomous agents; they struggle with very complicated tasks requiring deep domain expertise, long multi-step execution, nuanced reasoning, dynamic adaptation, or strict long-context consistency dependencies, as seen in scientific research Feng et al. (2024); Yehudai et al. (2025). (3) Safety and Ethical Risks escalate with autonomy; agents can take unintended harmful actions, amplify societal biases present in training data, or create accountability gaps, particularly in critical decision-making scenarios involving finance, healthcare, or security Mitchell et al. (2025); Deng et al. (2024); Wang et al. (2024c).
@@
Figure 1: Overview of LLM-based Human-Agent Systems (LAM-HAS). LLM-HAS are interactive frameworks where humans actively provide additional information, feedback, or control during interaction with an LLM-powered agent to enhance system performance, reliability, and safety. The system is composed of five core components: Environment & Profiling (including environment settings, and role definitions, goals, and agent capabilities such as planning and memory), Human Feedback (with varying types, timing, and granularity), Interaction Types (collaborative, competitive, cooperative, or mixed), Orchestration (task strategy and temporal synchronization), and Communication (information flow structure and mode).
@@
The persistence of these challenges suggests that full autonomy may be unsuitable for many real-world applications Mitchell et al. (2025); Natarajan et al. (2025) and underscores a crucial insight often overlooked in the drive for pure automation: the indispensable role of human involvement. Humans are frequently needed to provide additional information, essential clarification, or domain knowledge, offer vital feedback and corrections, and exercise necessary oversight and control. These motivate a paradigm shift towards systems explicitly designed for human-agent collaboration: LLM-based Human-Agent Systems (LLM-HAS).
@@
While surveys on LLM-based autonomous agents Wang et al. (2024a); Li et al. (2024a), multi-agent systems Tran et al. (2025); Wu et al. (2025), and specific applications exist Wang et al. (2025b); Peng et al. (2025), a dedicated synthesis focusing specifically on LLM-based human-agent systems is lacking. This survey fills the gap by providing a comprehensive and structured overview of the LLM-HAS. It clarifies the fundamental concepts and systematically presents its core components, emerging applications, and unique challenges and opportunities within this specific niche. To the best of our knowledge, this is still the first survey on LLM-based human-agent systems. We aim to consolidate current knowledge and inspire further research and application in this rapidly evolving interdisciplinary field. We maintain an open-source GitHub repository to provide a sustainable resource complementing our survey paper.
@@
This survey is organized as follows: Section 2 defines and formulates LLM-HAS. Section 3 details the core components shaping the human-agent systems (e.g., human feedback, interaction type, orchestration and communication protocols). Section 4 explores diverse application domains. Section 5 presents open-source implementation frameworks as well as datasets and benchmarks. Section 6 discusses key challenges and future opportunities.
@@
Type, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Evaluative, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ AgentCoord Pan et al. (2024a), CowPilot Huq et al. (2025),
@@
ConvCodeWorld Han et al. (2025), GDfC Wang et al. (2025c), DigiRL Bai et al. (2024), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Corrective, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Collaborative Gym Shao et al. (2024), ReHAC Feng et al. (2024),
@@
AI Chains Wu et al. (2022), HRC DMP Liu et al. (2024a), BPMN Ait et al. (2024), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Guidance, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ FineArena Xu et al. (2025), InteractGen Sun et al. (2024b),
ConvCodeWorld Han et al. (2025), DPT Agent Zhang et al. (2025), Organized Teams Guo et al. (2024b), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Implicit, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), AXIS Lu et al. (2024), AssistantX Sun et al. (2024a),
EmoAgent Qiu et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Feedback
Granularity, fill= HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Holistic, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Help Feedback Mehta et al. (2024), PARTNR Chang et al. (2024), A2C Tariq et al. (2025),
CoELA Zhang et al. (2024a), DPT Agent Zhang et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font =] ] [ Segment, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ CowPilot Huq et al. (2025), SWEET-RL Zhou et al. (2025), PDFChatAnnotator Tang et al. (2024),
TaPA Wu et al. (2023), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Feedback
Phase, fill= HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ Initial Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ AgentCoord Pan et al. (2024a), AssistantX Sun et al. (2024a), Hierarchical AgentLiu et al. (2023b), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ During Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ ConvCodeWorld Han et al. (2025), HRC Manipulation Liu et al. (2023a), A2C Tariq et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Post-Task, fill=HumanFeedbackLight, text width= 5.5em, anchor=center, align=center, font=[ GDfC Wang et al. (2025c), SOTOPIA Zhou et al. (2024),DigiRL Bai et al. (2024),
EmoAgent Qiu et al. (2025), leaf, fill=HumanFeedbackLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] ] [ Interaction
@@
(Sec 3.3), fill=interactClr, text width= 7em, anchor=center, align=center, font=[ Collaboration, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Supervision, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ EasyLAN Pan et al. (2024b), CowPilot Huq et al. (2025), Hierarchical Agent Liu et al. (2023b),
ConvCodeWorld Han et al. (2025), HRC Manipulation Liu et al. (2023a), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Delegation, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Collaborative Gym Shao et al. (2024), GDfC Wang et al. (2025c), HRC Manufa Lim et al. (2024),
PaLM-E Driess et al. (2023), DigiRL Bai et al. (2024), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Coordination, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ MTOM Zhang et al. (2024b), Ask Before Plan Zhang et al. (2024c), MindAgent  Gong et al. (2023),
MetaGPT  Hong et al. (2023), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] [ Cooperation, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ CoELA Zhang et al. (2024a), InteractGen Sun et al. (2024b), PARTNR Chang et al. (2024),
MAIH Wang et al. (2024c), leaf, fill=interactLight, text width=44em, anchor=center, align=center, text centered, font = ] ] ] [ Competition, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ SOTOPIA Zhou et al. (2024), leaf, fill=interactLight, text width=51.6em, anchor=center, align=center, text centered, font = ] ] [ Coopetition, fill=interactLight, text width= 6em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), SOTOPIA Zhou et al. (2024), leaf, fill=interactLight, text width=51.6em, anchor=center, align=center, text centered, font = ] ] ] [ Orchestration
@@
Help Feedback Mehta et al. (2024), Ask Before Plan Zhang et al. (2024c), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] [ Simultaneous, fill=orchestLight, text width= 6.5em, anchor=center, align=center, font=[ MTOM Zhang et al. (2024b), HRT-ML Liu et al. (2024b), DPT Agent Zhang et al. (2025),
@@
HRC Manufa Lim et al. (2024), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] [ Asynchronous, fill=orchestLight, text width= 6.5em, anchor=center, align=center, font=[ Collaborative Gym  Shao et al. (2024), MetaGPT Hong et al. (2023), TaPA Wu et al. (2023), leaf, fill=orchestLight, text width=42.2em, anchor=center, align=center, text centered, font = ] ] ] ] [ Communication
(Sec 3.5), fill=communClr, text width= 7em, anchor=center, align=center, font=[ Structure, fill=communLight, text width= 5em, anchor=center, align=center, font=[ Centralized, fill=communLight, text width= 6em, anchor=center, align=center, font=[ Drive As You Speack Cui et al. (2024), Organized Teams Guo et al. (2024b),
HRC Manufa Lim et al. (2024), DigiRL Bai et al. (2024), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Decentralized, fill=communLight, text width= 6em, anchor=center, align=center, font=[ Prison Dilemm Jiang et al. (2025), InteractGen Sun et al. (2024b), TaPA Wu et al. (2023),
MINT Wang et al. (2024b), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Hierarchical, fill=communLight, text width= 6em, anchor=center, align=center, font=[ FineArena Xu et al. (2025), AgentCoord  Pan et al. (2024a),
EasyLAN Pan et al. (2024b), Hierarchical Agent Liu et al. (2023b), EmoAgent Qiu et al. (2025), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] [ Mode, fill=communLight, text width= 5em, anchor=center, align=center, font=[ Conversation, fill=communLight, text width= 6em, anchor=center, align=center, font=[ DigiRL Bai et al. (2024), WebLINX Lù et al. (2024), Organized Teams Guo et al. (2024b),
ConvCodeWorld Han et al. (2025), Help Feedback Mehta et al. (2024), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Observation, fill=communLight, text width= 6em, anchor=center, align=center, font=[ AXIS Lu et al. (2024), Attentive Supp Tanneberg et al. (2024a), EasyLAN Pan et al. (2024b),
MineWorld Guo et al. (2025), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] [ Message
Pool, fill=communLight, text width= 6em, anchor=center, align=center, font=[ SMALL Wang et al. (2024c), HRT-ML Liu et al. (2024b), MetaGPT Hong et al. (2023), leaf, fill=communLight, text width=45em, anchor=center, align=center, text centered, font = ] ] ] ] ]
Figure 2: Taxonomy of LLM-based Human-Agent Systems. A more detailed and structured categorization of representative works is provided in the appendix (Table D.1 and D.2).